# Arrival Time Prediction for Autonomous Shuttle Services in the Real World: Evidence from Five Cities

### v1

@@ Abstract
Urban mobility is on the cusp of transformation with the emergence of shared, connected, and cooperative automated vehicles. Yet, for them to be accepted by customers, trust in their punctuality is vital. Many pilot initiatives operate without a fixed schedule, thus enhancing the importance of reliable arrival time (AT) predictions. This study presents an AT prediction system for autonomous shuttles, utilizing separate models for dwell and running time predictions, validated on real-world data from five cities. Alongside established methods such as XGBoost, we explore the benefits of integrating spatial data using graph neural networks (GNN). To accurately handle the case of a shuttle bypassing a stop, we propose a hierarchical model combining a random forest classifier and a GNN. The results for the final AT prediction are promising, showing low errors even when predicting several stops ahead. Yet, no single model emerges as universally superior, and we provide insights into the characteristics of pilot sites that influence the model selection process. Finally, we identify dwell time prediction as the key determinant in overall AT prediction accuracy when autonomous shuttles are deployed in low-traffic areas or under regulatory speed limits. This research provides insights into the current state of autonomous public transport prediction models and paves the way for more data-informed decision-making as the field advances.
Index Terms:
Machine Learning, Autonomous Public Transport, Arrival Time Prediction, Graph Neural Networks
@@ I Introduction
Shared, connected, and cooperative automated vehicles offer a unique opportunity for a fundamental change in urban mobility. They can provide seamless door-to-door mobility of people and freight delivery services, which can lead to more accessible, greener, and more sustainable cities - provided they are integrated into an effective public transport system. For users to accept this new mode of transportation, they need to trust its punctuality, making reliable AT prediction systems crucial for a satisfactory user experience. Fortunately, connected automated vehicles also constitute a unique Big Data source. When coupled with the latest developments in Machine Learning and Artificial Intelligence, this data can enable the implementation of AT prediction systems. But the question arises - Do the findings about established machine learning models for conventional busses transfer to their automated counterparts? With the recent developments in both technological and regulatory matters, autonomous public transport is still in pilot testing and is constrained by local regulations. The data that these pilot projects yield, though invaluable, is limited and varies in quality. Thus, it is unclear if this data can serve as a foundation for reliable prediction models. Moreover, so far, the primary function of autonomous shuttles in real-life case studies has been to connect existing public transport trunk lines with remote areas, such as schools, residential care facilities, and shopping centers. Typically, there is no historical data, such as passenger counts, from existing infrastructure that can be used as an indication for this new service. Furthermore, these shuttles operate without fixed schedules, rendering schedule adherence infeasible and thus making AT prediction especially challenging, yet, at the same time, an even more important task.
To our knowledge, this paper is the first to evaluate various AT prediction models incorporating dwell time and running time predictions in the context of real-life case studies of autonomous shuttles. We approach the AT prediction problem as two distinct regression problems for dwell and running time and put a spectrum of machine learning models, from the established tree-based methods to more complex graph neural networks, to the test, all while drawing insights from data from five diverse pilot sites. This work is an initial study of the performance of these models in autonomous public transport within a real-world autonomous public transport setting.
@@
The structure of the paper is as follows. First, we give an overview of related work on bus AT prediction. Following, we introduce our methodology on our segment-based approach to distinguish between dwell and running time along with the corresponding prediction models. Then, we discuss the data pre-processing steps before presenting the final results. Finally, we conclude our work while giving future research directions.
@@ II-A Bus Arrival Time prediction
Bus AT prediction is a large research area within Intelligent Transport Systems. Historically, Kalman Filters has been a common approach to AT predictions due to their capability to maintain states between predictions and filter noise [1, 2, 3]. More recently, Machine Learning models have shown good performance for AT predictions. Shallow Machine Learning techniques like Support Vector machines [4] and Random Forest [5] have been utilized for AT; however, most work has focused on models based on Artificial Neural Networks (ANN), e.g. [6] created a hierarchical ANN model and showed that it outperformed Kalman Filters for short-distance routes. Long Short-Term Memory models (LSTMs) have been popular for predicting AT due to their ability to capture the non-linear behavior of time-series [7, 8, 9]. In [10], the authors combine LSTMs with Convolutional Neural Nets to learn the temporal and spatial correlations between bus stops. [11] used Graph Neural Networks (GNNs) for city-wide bus travel time estimation. However, unlike our approach, they do not split travel time into dwell time and running time.
@@
In [12], the authors show that dwell time is among the most important factors determining AT. As such, it can be beneficial for AT prediction to split the task into predicting the dwell and running time separately and then accumulate the predictions afterward. In [13], the authors propose a segment-based approach to predict bus AT by dividing bus routes into dwelling and transit segments. The authors find bus stops by clustering GPS signals and extracting dwell and running times from the GPS traces. The final AT predictions are then generated by accumulating the independent dwell and running time predictions. [10] also argue for splitting bus AT models into travel and dwell time models. The authors utilize LSTM and CNNs for the running time predictions and a more simple exponential smoothing model for dwell time predictions. For a more in-depth survey of general AT prediction research, we refer to [14].
@@
As autonomous shuttles are a new technology, they have not yet seen much real-world application, and as such, the research on predicting arrival times for automated shuttles is lacking. [15] did an initial study of the running time of autonomous shuttles. The authors used the current location along the current segment, the elapsed time on the current segment, and the shuttle’s speed and acceleration as input to a Gradient Boosting Regression Tree model to predict the remaining time on the current segment. The authors added artificial stops to the route as the study was done in the early stages of an autonomous shuttle pilot.
As our work is based on data from more advanced pilot sites, we are, to the best of the authors’ knowledge, the first to do dwell and running time predictions on data from autonomous shuttles in real-life deployment. Furthermore, we are also the first to utilize GNNs and hierarchical GNNs for split dwell and running time predictions.
@@ III Methods
Dwell times are mainly impacted by the uncertain numbers of boarding passengers and hence show structural differences to running times, which are primarily impacted by traffic conditions. Therefore, in this work, we adopt a segment-based approach, employing two separate models to predict the running and dwell times. Figure 1 provides an overview of the segment-based approach. The dwell time corresponds to the time the shuttle is stationary at a shuttle stop and the running time to the time it travels between two stops including acceleration and deceleration time. To obtain the final travel time $T_{i,j}$ of the autonomous shuttle from stop $i$ to stop $j$, we aggregate the predictions for the individual segments for the dwell time $dt$ and running time $tt$ along the route:
$T_{i,j}=\sum_{k=i}^{j-1}tt_{k,k+1}+\sum_{k=i+1}^{j-1}dt_{k}.$ (1)
@@
IV Dwell and running time prediction models
Given that current data on autonomous shuttles is confined to test pilots, our models must operate on limited input. Many studies have equipped their models with real-time traffic information from available car data i.e. taxi data [13] to obtain accurate running time predictions. However, as autonomous shuttles often operate on both shared public roads and private lanes or share lanes only with pedestrians and cyclists, leveraging standard road traffic speed estimates becomes unfeasible. Moreover, these estimates might not be as impactful in the case of autonomous shuttles, as they adhere to specific speed regulations designed for such vehicles, which are often lower than the road’s general speed limit. Hence, our approach adopts a time-series-based approach, where we include previous observations of the running and dwell time, respectively, on the segment into the regression task. However, given the data sparsity resulting from pilot testing, the most recent observations can quickly become outdated and, thus, less reliable. Hence, our models only include the two previous observations per segment. As such, our utilized features include the current position of the shuttle, current weather data, current time, and the ID of the segment and the vehicle in question, along with the two most recent observations. So given a vehicle id $i$, segment id $j$ and time $t$ the input to our prediction model are:
$\boldsymbol{x}_{i,j,t}=\left[\tau_{t},w_{t},v_{i},s_{j},l^{(1)}_{j,t},l^{(2)}_{j,t}\right]$ (2)
where $\tau_{t}$ is the geometrically transformed time-of-day and day-of-week features, $w_{t}$ is the temperature, precipitation, and windspeed, $v_{i}$ is a one-hot encoding of the vehicle id, $s_{j}$ is a one-hot encoding of the segment id and $l^{(1)}_{j,t}$ and $l^{(2)}_{j,t}$ are the most recent and second most recent lag of segment $j$ at time $t$. The lags are either historical dwell time or running time, depending on which is being predicted. Then we define a model $\mathcal{F}$ that takes in the input data $x_{i,j,t}$ and predicts the next running or dwell time for that vehicle on the segment, i.e.
$\boldsymbol{x}_{i,j,t}\xrightarrow{\mathcal{F}}y_{i,j,t}.$ (3)
Our prediction models employ machine learning techniques that have proven successful in predicting both the dwell and the running times of conventional buses. Studies highlighted the competitive performance of tree-based methods, especially XGBoost, in both dwell time [16, 17], travel time ([18] and running time predictions [19]). XGBoost [20], which stands for eXtreme Gradient Boosting, is an optimized gradient boosting algorithm that builds on decision trees. In addition, our work also examines three deep learning models:
  • A Multi-Layer Perceptron (MLP) that consists of fully connected layers.
  • A Graph Convolutional Network (GCN) [21]. GCNs share information among nodes in the graph utilizing a permutation-invariant propagation rule. This way GCNs are able to capture spatial correlations in the data.
  • A hierarchical model that combines a Random Forest (RF) [22] followed by GCN. Drawing from the tools from zero-inflated regression, the binary RF classifier handles zero observation while the downstream GCN regresses the non-zero observations. We call this model RF-GCN.
Given the high temporal resolution of our data sets, when the shuttle is skipping a stop, we can determine a dwell time of exactly 0s. In pilot tests, the demand is very volatile, resulting in a data set with a potentially significant number of zeros. Detecting those zeros is important for AT prediction, as overestimating the dwell time when it is actually zero leads to an overestimation of the total travel time. As a result, the shuttle departs earlier than predicted, which significantly deteriorates the customer experience. Further, for GCN-based approaches, the input is a matrix $\boldsymbol{X}$ where the $k$’th row is $x_{i,k,t}$. The GCN model then predicts a vector, $\boldsymbol{y}_{i,t}$, with a prediction for each node. The prediction for segment $j$ can then be taken out as the $j$’th position.
V Data
The data utilized for our models were gathered from five cities conducting pilot tests with autonomous shuttles, all of which are part of the SHOW (SHared automation Operating models for Worldwide adoption) project [23]. The SHOW Project is a large-scale EU initiative aiming to demonstrate the viability and effectiveness of shared, connected, cooperative, and electrified fleets of Autonomous Vehicles (AVs) in real-world urban demonstrations across 20 cities in Europe. Altogether, the project will deploy a fleet of more than 79 SAE L4/L5 AVs for both passenger and cargo transport in dedicated lanes as well as mixed traffic. In this paper, we test our models on data from five of these pilot sites (Table I), where tests have been ongoing for several months. Currently, all vehicles are automated at level 4. The pilot sites in Table I represent not only different stages of deployment but also some have a particular focus, such as teleoperation or the challenge of interfacing with different road users.
The quality of the provided data varies significantly with the pilot site; while all provide GPS positioning and speed data, some have deviations. For instance, the speed data from Linköping is noisy, and in Les Mureaux, the sampling rate of the GPS data is decreased mid-pilot. Such variations are not limited to our study but are representative of many other pilots, both current and upcoming. Linköping stands out as the site with the longest ongoing pilot, as well as the longest route. At the other end is Graz, with the shortest route and a very limited number of observations. These distinct attributes across pilot sites provide an optimal foundation for our model evaluations.
TABLE I: Data Stats
Site Country #Routes #Stops Dates Avg. speed Max speed #Vehicles #Observations
Linkoping Sweden 2 15 12/21-06/23 07.7 km/h 22 km/h 3 48 679
Tampere Finland 1 07 12/21-03/22 08.2 km/h 64 km/h 2 06 696
Les Mureaux France 3 07 11/22-02/23 02.2 km/h 06 km/h 2 04 580
Madrid Spain 1 05 06/22-02/23 04.1 km/h 36 km/h 2 07 096
Graz Austria 2 04 10/22-06/23 14.3 km/h 57 km/h 1 00 359
V-A Data pre-processing
To enable a clear distinction between run and dwell time in public transport systems, careful data pre-processing is crucial. We consider two scenarios: One where we have both speed and GPS data and another where we rely solely on GPS positioning. In the first scenario, with the combination of speed and GPS data, we have explicit information about when a shuttle is in motion or stationary. To determine the dwell time, we define a shuttle as dwelling when it is at a standstill (speed of 0 km/h) and within a predetermined radius of a stop. As soon as the shuttle starts moving, we start the running time calculation for the corresponding route segment. If the shuttle bypasses a stop without stopping, we consider the dwell time as zero.
However, in some cases, e.g. in Linköping, GPS positioning might be the only data at the disposal, and the information about the shuttle’s movement is less explicit. In this scenario, we establish a threshold for GPS differences to detect when a shuttle dwells at a stop while accounting for the noise (GPS jitter) in the GPS data. Dwell times are then calculated based on periods of minimal movement, as defined by the threshold. A predefined stop radius remains equally relevant here to identify when the shuttle is at a stop.
There are additional cases that require specific attention in our pre-processing. There are instances when a shuttle is parked between runs or traveling to and from its depot without turning off the GPS device. We manually identify these relevant locations and exclude instances where the shuttle deviates from its designated route from our data set. We also handle situations where two stops on opposite sides of the road have overlapping radii by hard-coding all possible stop orders in the pre-processing to ensure precise calculations. This way, we can precisely determine the relevant shuttle stop for each data point, eliminating ambiguity and potential errors. Other site-dependent adjustments must be considered, such as the change in GPS sampling frequency in Les Mureaux. Here, the GPS sampling rate was reduced mid-pilot, affecting time calculations, which required us to discard the impacted data for consistency. For the GNN, the data must be structured as a graph with nodes and edges linking these nodes based on an adjacency matrix. In the context of dwell time prediction, each node represents a stop, while for the running time prediction, the nodes correspond to the segments between stops. The connectivity of the graph follows the different routes of the shuttle. We illustrate the two graphs for Linköping in Figure 2.

### v2

# A Large-scale Analysis on the Use of Arrival Time Prediction for Automated Shuttle Services in the Real World

@@ Abstract
Urban mobility is on the cusp of transformation with the emergence of shared, connected, and cooperative automated vehicles. Yet, for them to be accepted by customers, trust in their punctuality is vital. Many pilot initiatives operate without a fixed schedule, enhancing the importance of reliable arrival time (AT) predictions. This study presents an AT prediction system for automated shuttles, utilizing separate models for dwell and running time predictions, validated on real-world data from six cities. Alongside established methods such as XGBoost, we explore the benefits of leveraging spatial correlations using graph neural networks (GNN). To accurately handle the case of a shuttle bypassing a stop, we propose a hierarchical model combining a random forest classifier and a GNN. The results for the final AT prediction are promising, showing low errors even when predicting several stops ahead. Yet, no single model emerges as universally superior, and we provide insights into the characteristics of pilot sites that influence the model selection process and prediction performance. Finally, we identify dwell time prediction as the key determinant in overall AT prediction accuracy when automated shuttles are deployed in low-traffic areas or under regulatory speed limits. Our meta-analysis across six pilot sites in different cities provides insights into the current state of autonomous public transport prediction models and paves the way for more data-informed decision-making as the field advances.
@@ I Introduction
Shared, connected, and cooperative autonomous vehicles offer a unique opportunity for a fundamental change in urban mobility. They can provide seamless door-to-door mobility of people and freight delivery services, which can lead to more accessible, greener, and more sustainable cities - provided they are integrated into an effective public transport system. For users to accept this new mode of transportation, they need to trust its punctuality, making reliable AT prediction systems crucial for a satisfactory user experience. Fortunately, connected autonomous vehicles also constitute a unique Big Data source. When coupled with the latest developments in Machine Learning and Artificial Intelligence, this data can enable the implementation of AT prediction systems. However, an important question arises: do the findings about established machine learning models for conventional buses transfer to their automated counterparts? With the recent developments in both technological and regulatory matters, autonomous public transport is still in pilot testing and is constrained by local regulations. The data that these pilot projects yield, though invaluable, is limited and varies in quality. Thus, it is unclear if this data can serve as a foundation for reliable prediction models. Moreover, so far, the primary function of automated shuttles in real-life case studies has been to connect existing public transport trunk lines with remote areas, such as schools, residential care facilities, and shopping centers. Typically, there is no historical data, such as passenger counts, from existing infrastructure that can be used as an indication for this new service. Furthermore, these shuttles often operate without fixed schedules, rendering schedule adherence infeasible and thus making AT prediction especially challenging, yet, at the same time, an even more important task.
To our knowledge, this paper is the first to systematically benchmark a broad spectrum of AT prediction models for automated shuttles, while explicitly separating and modeling running-time and dwell-time components. Because automated, smaller shuttle services are often deployed to serve previously remote areas, demand is highly volatile, and thus dwell-time data are heavily zero-inflated: stops are often skipped when no passengers are present. To handle this, we introduce a hierarchical framework of a Random-Forest classifier and a Graph Neural Network regressor. Further, this work is, to our knowledge, the first use of GNNs for dwell-time prediction, exploiting spatial correlations between stops, whereas earlier studies rely on simpler models [Rashidi2013ApproximationAS, Chen2007UsingAP, MA2019536, PETERSEN]. We evaluate all models on real-world data from six heterogeneous pilot sites, spanning different vehicle types, routes, and operating contexts, thereby providing insights that are valuable to both researchers and practitioners working with emerging automated-shuttle systems.
@@
First, we review related work on bus AT prediction. Next, we present our segment-based approach for separating dwell and running time, along with the corresponding prediction models. We then describe the data pre-processing steps before presenting the final results. Finally, we conclude and outline future research directions.
@@ II Related work
Bus AT prediction is a large research area within Intelligent Transport Systems. Historically, Kalman Filters have been a common approach to AT predictions due to their ability to maintain states between predictions and filter out noise [Chen2004ADB, Wall1998AnAF, Sinn2012PredictingAT, Achar2020Kalman]. More recently, Machine Learning (ML) models have demonstrated promising performance for AT prediction tasks. Shallow ML techniques such as Support Vector machines [Yu2006BusAT] and Random Forests [Li2017BusAT] have been explored; however, most work has focused on models based on Artificial Neural Networks (ANN). For instance, [Lin2013RealTimeBA] developed a hierarchical ANN model, which significantly outperformed Kalman Filters for short-distance routes. Kalman filters remain attractive for their recursive updates and explicit uncertainty measurement, but recent studies show that deep ML models (LSTM, XGBoost) and, especially, GNNs offer benefits on bus AT prediction by exploiting high-dimensional context and network-wide spatial correlations, with hybrid ML + KF schemes now emerging as a promising combination [Bai, Achar, Shen] Long Short-Term Memory models (LSTMs) have become popular for AT prediction due to their ability to capture the non-linear behavior of time-series data, effectively modeling temporal dependencies [Agafonov2019BusAT, Pang2019LearningTP, Lingqiu2019ALB]. In [PARSLOV2023], the authors integrate uncertainty estimation into recurrent neural networks to predict running times, improving decision-making for transfer synchronization. In [PETERSEN], the authors combine LSTMs with Convolutional Neural Networks to capture both the temporal and spatial correlations between bus stops. [Ma2022MultiattentionGN] uses Graph Neural Networks (GNNs) for city-wide bus travel time estimation. However, unlike our approach, their model does not separate travel time into dwell and running time.
@@
In [Chen2007UsingAP], the authors, using simple ANNs, identify dwell time as among the most important factors determining AT using simple ANNs. As such, it can be beneficial to split the task into predicting the dwell and running time separately. In [MA2019536], the authors propose a segment-based approach to predict bus AT by dividing bus routes into dwelling and transit segments. The final AT predictions obtained by shallow ML approaches or simple ANNs are then generated by accumulating the independent predictions. Similarly, [PETERSEN] advocate for splitting bus AT models into travel and dwell time models, utilizing LSTM and CNNs for the running time predictions and a simpler exponential smoothing model for dwell time predictions. For a more in-depth survey of general AT prediction research, we refer to [Singh2022ARO].
@@
As automated shuttles are a new technology with limited real-world deployment, the research on predicting AT for automated shuttles is lacking. [Antypas] did an initial study of the running time of automated shuttles. They used factors such as location, elapsed time, speed, and acceleration as inputs for a Gradient Boosting Regression Tree model to predict the remaining segment time. The authors added artificial stops to the route as the study was done in the early stages of an automated shuttle pilot.
Building on data from advanced pilot sites, our work, to the best of the authors’ knowledge, is the first to benchmark a range of prediction models on separate dwell and running time prediction tasks with data from automated shuttles in real-world deployments. Moreover, our approach is the first to use GNNs and hierarchical GNN architectures to separately predict dwell and running times, incorporating spatial information into dwell time prediction - a component overlooked in previous research.
@@ III Methods
Dwell times are mainly impacted by the uncertain number of boarding passengers and hence show structural differences to running times, which are primarily impacted by traffic conditions. Therefore, in this work, we adopt a segment-based approach, employing two separate models to predict the running and dwell times. Figure 1 provides an overview of the segment-based approach. The dwell time corresponds to the time the shuttle is stationary at a shuttle stop, and the running time to the time it travels between two stops, including acceleration and deceleration time. To obtain the final travel time $T_{i,j}$ of the automated shuttle from stop $i$ to stop $j$, we aggregate the predictions for the individual segments for the dwell time $dt$ and running time $tt$ along the route:
$T_{i,j}=\sum_{k=i}^{j-1}tt_{k,k+1}+\sum_{k=i+1}^{j-1}dt_{k}.$
@@
III-A Dwell and running time prediction models
Given that current data on automated shuttles is confined to test pilots, our models must operate on limited input. Many studies have equipped their models with real-time traffic information from available car data, e.g., taxi data [MA2019536], to obtain accurate running time predictions. However, as automated shuttles often operate on both shared public roads and private lanes or share lanes only with pedestrians and cyclists, leveraging standard road traffic speed estimates becomes unfeasible. Moreover, these estimates might not be as impactful in the case of automated shuttles, as they adhere to specific speed regulations designed for such vehicles, which are often lower than the road’s general speed limit. Hence, our approach adopts a time-series-based approach, where we include previous observations of the running and dwell time, respectively, on the segment into the regression task. However, given the data sparsity resulting from pilot testing, the most recent observations can quickly become outdated and, thus, less reliable. Hence, our models only include the two previous observations per segment. As such, our features include the current position of the shuttle, current weather data, current time, and the ID of the segment and the ID of the vehicle in question (since vehicles can have different manufacturers and therefore different behaviors), along with the two most recent observations. Given vehicle id $i$, segment id $j$ and time $t$, the input to our prediction models is:
$\boldsymbol{x}_{i,j,t}=\left[\tau_{t},w_{t},v_{i},s_{j},p_{j},l^{(1)}_{j,t},l^{(2)}_{j,t}\right$],
where $\tau_{t}$ are the geometrically transformed time-of-day and day-of-week features, $w_{t}$ is a vector representing the temperature, precipitation, and wind speed at an hourly resolution, $v_{i}$ is a one-hot encoding of the vehicle id, $s_{j}$ is a one-hot encoding of the segment id, $p_{j}$ the position of the stop or start of running time segment (latitude and longitude), and $l^{(1)}_{j,t}$ and $l^{(2)}_{j,t}$ are the most recent and second most recent lag of segment $j$ at time $t$. The lags are either historical dwell time or running time, depending on which is being predicted. Then, we define a model $\mathcal{F}$ that takes in the input data $x_{i,j,t}$ and predicts the next running or dwell time for that vehicle on the segment,
$\boldsymbol{x}_{i,j,t}\xrightarrow{\mathcal{F}}\hat{y}_{i,j,t}.$
Specifically, $\hat{y}_{i,j,t}$ corresponds to the predicted running time on segment $j$ given entry time at $t$ or dwelling time at stop $j$ given arrival at the stop $j$ at time $t$. In graph neural network-based approaches, where the input must be structured as a graph with nodes and their respective features, the input becomes a matrix $\boldsymbol{X}$ where the $k$’th row is $x_{i,k,t}$. The GCN model then predicts a vector, $\boldsymbol{y}_{i,t}$, with a prediction for each node. The prediction for segment $j$ can then be taken out as the $j$’th position.
Specifically, for the GNN, the input data must be structured as a graph with nodes and edges linking these nodes based on an adjacency matrix. For the dwell time model, each node represents a stop. For the running-time model, each node represents the road segment between two consecutive stops. The connectivity of the input graph corresponds to the physical shuttle routes. Two nodes are connected if, in at least one of the shuttle routes, they are visited consecutively. We illustrate the graphs for Linköping in Figure 2.
Our feature set is designed to allow temporal generalization, using temporal and weather information. This approach addresses the most common scenario that pilot sites aim to evaluate. For cross-city transferability, we anticipate that either initial fine-tuning on a dataset from the new city or an extension of the current feature set, with e.g., demand estimates, POI densities, or road types, to be necessary.
To evaluate the different prediction models, we compare their performance against two naive baseline models:
  • a Lag model, which predicts the most recent observation on the respective segment;
  • a Mean regressor, also often referred to as a historical average model, which predicts the mean of the respective segment.
We employ machine learning techniques that have proven successful in predicting both the dwell and the running times of conventional buses. Studies highlighted the competitive performance of tree-based methods, especially XGBoost, in both dwell time [10125868, Rashidi_dwell], travel time [travel_time_xgboost], and running time predictions [travel_time_xgboost2]. We include the following tree-based models:
  • XGBoost [xgboost], which stands for eXtreme Gradient Boosting, is an optimized gradient boosting algorithm that constructs decision trees sequentially, with each new tree aiming at minimizing the residual errors of the previous ones.
  • Random Forest [RF], a simpler yet highly effective tree-based model. Random Forest is an ensemble method that builds multiple decision trees and aggregates their results to improve prediction accuracy and robustness.
TABLE I: Pilot site characteristics.
Site Country #Routes #Stops #Vehicles Shuttle types (#types, #passengers) Speed limit Deployment
Linköping Sweden 2 15 3 Minibus (2, 11/13) 30-40km/h Campus+Residental
Tampere Finland 1 07 2 Van (1, 7) 40km/h Sub-urban+Residential
Les Mureaux France 3 07 2 Minibus (1, 12) 30-50km/h Industrial area
Madrid Spain 1 05 2 Minibus (2, 30/105) NA Bus depot
Graz Austria 2 04 1 Passenger car 30km/h Retail area
Carinthia Austria 2 08 1 Minibus (1, 11) 30-50km/h Residental areal
In addition, we examine three deep learning models:
  • A Multi-Layer Perceptron (MLP) that consists of fully connected layers, capturing non-linear relationships in the data through layer-wise transformations. It provides a baseline deep-learning approach for non-graph inputs and has been traditionally applied to both tasks [Rashidi_dwell, GURMU201445].
  • A Graph Convolutional Network (GCN) [kipf2017semi], which propagates information among nodes in a graph using a permutation-invariant propagation rule. This enables GCNs to capture spatial dependencies within the data, making them well-suited for tasks with spatially correlated features. As a result, GCNs have proven successful in numerous transportation applications, such as traffic flow prediction [Li2017DiffusionCR], and fit our prediction tasks as both inherently include spatial correlations.
  • A hierarchical model that combines a Random Forest classifier with a GCN regressor (RF-GCN). In smaller shuttle services, demand volatility yields a dataset with a potentially high number of zero dwell times. Accurately detecting these zeros is essential for AT prediction, as overestimating the dwell time leads to an overestimation of the total travel time. To address this, we propose the RF-GCN model. Drawing from the tools of zero-inflated regression, the binary RF classifier is used to detect zero observations, while the downstream GCN regresses the non-zero observations. Concretely, the RF classifies whether the shuttle will stop vs. skip using binary labels ($z_{i}=1,\text{if }y_{i}>0$ and $z_{i}=0,\text{if }y_{i}=0$). Accordingly, it is trained on a modified classification dataset $D_{\text{cls}}\;=\;\{(x_{i},z_{i}):i=1,\dots,N\}$. The GNN regressor predicts the duration of the dwell event $\tilde{y}_{i}$ conditional on stopping. It is trained on a subset of the original dataset, only including non-zero targets $D_{\text{reg}}\;=\;\{(x_{i},y_{i}):y_{i}>0\}$. Excluding zeros focuses the regressor on actual boarding and alighting durations and avoids bias from zero-inflation. Crucially, both stages are trained independently on their respective datasets. At inference, the classifier outputs the probability of a stop being served; if the predicted class is $\hat{z}_{i}=0$ (”skip”), we set $\hat{y}_{i}=0$. If $z_{i}=1$ (”stop”), we query the GCN for a continuous estimate of the dwell time $\tilde{y}_{i}$ and set $\hat{y}_{i}=\tilde{y}_{i}$.
We conduct grid searches (RF, XGBoost) and random searches (MLP, GCN) on validation sets to optimize hyperparameters. For simpler models (RF, XGBoost), we select hyperparameters for each pilot site. For the deep learning models, due to higher computational time, we use the same parameters across sites. The hyperparameters and training scripts for all evaluated models are available.  [^1] https://github.com/carolinssc/Arrival-time-prediction-autonomous-shuttles
TABLE II: Summary data statistics for each pilot site. #Observation indicates running/ dwell times. Running and dwell time stats present mean values and coefficient of variation in seconds. For dwell time, the fraction of zero observations is additionally provided in parentheses.
Site Dates Avg. speed Max speed #Observations #Passengers Running time stats Dwell time stats
Linköping 12/21-02/23 07.6 km/h 22 km/h 45 059/48 679 17 683 120, 0.14 28.0, 0.38, (0.06)
Tampere 12/21-03/22 08.5 km/h 64 km/h 06 177/06 696 08 643 091, 0.17 17.4, 1.61, (0.44)
Les Mureaux 11/22-02/23 02.9 km/h 06 km/h 04 580/04 847 01 084 174, 0.13 31.3, 0.41, (0.17)
Madrid 06/22-02/23 03.0 km/h 17 km/h 06 166/07 096 06 363 063, 0.34 31.2, 2.70, (0.75)
Graz 10/22-06/23 04.5 km/h 17 km/h $0$ 662/$0$ 359 00520 132, 0.18 15.3, 1.05, (0.17)
Carinthia 06/22-12/22 08.1 km/h 18 km/h 01 645/01 643  [^0] NA 146, 0.16 24.0, 1.00, (0.26)
IV Data
The data were gathered from six cities conducting pilot tests with automated shuttles, all of which are part of the SHOW (SHared automation Operating models for Worldwide adoption) project [SHOWweb]. The SHOW Project is a large-scale EU initiative aiming to demonstrate the viability and effectiveness of shared, connected, cooperative, and electrified automated vehicles (AVs) in real-world urban demonstrations across 20 cities in Europe. Altogether, the project deploys a fleet of more than 80 SAE L4/L5 AVs for both passenger and cargo transport in dedicated lanes as well as mixed traffic. In this paper, we test our models on data from six of these pilot sites (Tables I), where tests have been ongoing for several months. Currently, all vehicles are automated at level 4. The pilot sites in Table I represent not only different stages of deployment but also some have a particular focus, such as teleoperation or the challenge of interfacing with different road users. The variability in sample sizes, induced by the different pilot lengths and data collections, allows us to analyze the impact of sample size and operating conditions on the predictive performance of the different models. In particular, limited sample sizes are something that real-world pilot projects introducing new services inevitably face in practice, and the datasets used in this study align our research with real-world conditions.
The quality of the provided data varies with the pilot site; while all provide GPS positioning and speed data, some have deviations. For instance, the speed data from Linköping is noisy, and in Les Mureaux, the sampling rate of the GPS data is decreased mid-pilot. Such inconsistencies are not unique to our study but are representative of many other real-world pilots, both current and upcoming. As shown in Tables I and II, Linköping stands out as the site with the longest ongoing pilot, as well as the longest route, contrasting with Graz, which has the shortest route and a very limited number of observations and passengers. The majority of pilots deploy minibusses, with the exceptions of Tampere and Graz, which use vans and passenger cars, respectively. This difference is reflected in the mean dwell time, as these two sites exhibit roughly half the dwell time compared to the other pilots. While most pilots operate in suburban and residential areas, Madrid and Les Mureaux stand out. Madrid’s pilot, conducted within a private bus depot and deploying higher capacity vehicles, focuses on automated depot management and auto parking, thus yielding the highest variability in dwell and running times. On the other hand, Les Mureaux, operating for employees on a private company site, shows the second-lowest passenger count and the lowest variability in running time. These distinct pilot site characteristics provide an optimal foundation for our model evaluations since they provide a comprehensive representation of the broad array of characteristics that can be found in real-world deployments, thus allowing for an insightful meta-analysis of the results.
IV-A Data pre-processing
To enable a clear distinction between running and dwell time in public transport systems, careful data pre-processing is crucial. We provide an overview of the data preprocessing algorithm in Algorithm 1.
Algorithm 1 Travel and Dwell Time Calculation
1:Input: GPS data with coordinates, timestamps, and optionally speed; Shuttle routes
2:Output: Running times and dwell times between stops
3:procedure Calculate Times(data, stops)
4:  1. Assign stop IDs to GPS points based on proximity to the next valid stops in the route.
5:  2. Mark observation as stationary if speed <0.01<0.01 or GPS differences below jitter threshold.
6:  3. For each stop, compute dwell time as the total time the shuttle is stationary.
7:  4. For each pair of consecutive stops, compute running time as the time between departure and arrival.
8:end procedure
We consider two scenarios: One where we have speed and GPS data, and another where we rely solely on the latter. In the first scenario, with the combination of speed and GPS data, we have explicit information about when a shuttle is in motion or stationary. To determine the dwell time, we define a shuttle as dwelling when it is at a standstill (speed of 0 km/h) and within a predetermined radius of a stop. As soon as the shuttle starts moving, we start the running time calculation for the corresponding route segment. If the shuttle bypasses a stop without stopping, we consider the dwell time as zero.
However, in some cases, e.g., in Linköping, GPS positioning might be the only data available, and the information about the shuttle’s movement is less explicit. In this scenario, we establish a threshold for GPS differences to detect when a shuttle dwells at a stop while accounting for the noise (GPS jitter) in the GPS data. Dwell times are then calculated based on periods of minimal movement, as defined by the threshold. A predefined stop radius remains equally relevant here to identify when the shuttle is at a stop.
There are additional cases that require specific attention in our pre-processing. There are instances when a shuttle is parked between runs or traveling to and from its depot without turning off the GPS device. We manually identify these relevant locations and exclude instances where the shuttle deviates from its designated route from our data set. We also handle situations where two stops on opposite sides of the road have overlapping radii by hard-coding all possible stop orders in the pre-processing. This way, we can precisely determine the relevant shuttle stop for each data point, eliminating ambiguity and potential errors. Other site-dependent adjustments must be considered, such as the change in GPS sampling frequency in Les Mureaux. Here, the GPS sampling rate was reduced mid-pilot, affecting time calculations, which requires us to discard the impacted data for consistency. These aspects are not unique to our case study but represent challenges commonly encountered in real-world deployments. We emphasize the importance of addressing them to ensure the most accurate data preprocessing possible.